OpenStudy (anonymous):

the third term of an arithmetic sequence is 21, the eighth term is 56. the first term is….

OpenStudy (cwrw238):

if a is the first term we have 3rd term = a + 2d = 21 8th term = a + 7d = 56 solve this system of equations to find a and d

OpenStudy (cwrw238):

remember that nth term = a + (n-1)d where d = common difference, a = first term

OpenStudy (anonymous):

where does the 2 and 7 come from?

OpenStudy (cwrw238):

n-1 in a + (n-1)*d for 3rd term, n = 3 so n-1 = 2

OpenStudy (anonymous):

okay so was the first term 7?